OpenStudy (anonymous):

A number is four times another number. When the numbers are added, their sum is 40. What are the numbers.

OpenStudy (anonymous):

8 and 32

OpenStudy (anonymous):

lol, dhat will explain it ^_^

OpenStudy (anonymous):

I thought it was wanting me to turn it into a equations. Idk what I was doing.

OpenStudy (anonymous):

lol yes you can

OpenStudy (anonymous):

lets say the numbers are x and y. suppose y is 4 times x, then y = 4x their sum is 40 x+y = 40, so x+4x = 40, so x = 40/5 = 8

OpenStudy (anonymous):

a number (x) is (=) 4 times (x) another number(y) x = 4y ^_^, jut like dhat said

OpenStudy (anonymous):

lol, i flipped them, but same scenario :)

OpenStudy (anonymous):

lol yeah. tag! you're it!

OpenStudy (anonymous):

did you get it Meghan? ^_^

OpenStudy (anonymous):

LOL! it's been a while since I have last played tag

OpenStudy (anonymous):

heh. but seriously , Meghan, did you understand what happened there?

OpenStudy (anonymous):

give her a moment lol =P

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

good ^_^

OpenStudy (anonymous):

alright, good.